OpenStudy (istim):

Which pairs of planes are parallel and distinct and which are conincident? 2x+4y-7z-2=0 4x+6y-14z-8=0

OpenStudy (istim):

I tried looking through my notes and textbook for this one, but unless there is a direct connection, I don't understand how to apply it to this question.

OpenStudy (istim):

I managed to answer some unrelated questions beforehand, but I don't know how I managed to solve them...

OpenStudy (istim):

So I figured out the normals were [2,3,7] and [4,6,-14]

OpenStudy (istim):

I can see that I just have to multiply 2,3,7 by 2 to get the other plane. So would that be conidental or parallel and distinct?

OpenStudy (bahrom7893):

that should be parallel and distinct.. if u can just multiply them by a scalar, but i suck at this, so idk.. @Mertsj @Zarkon

OpenStudy (bahrom7893):

@amistre64 @satellite73 @myininaya

OpenStudy (bahrom7893):

@LagrangeSon678

OpenStudy (istim):

Wait. Maybe I'll try something out here.

OpenStudy (zarkon):

how is [2,3,7] the normal vector?

OpenStudy (istim):

Er, I just looked at my notes and inferred that...

OpenStudy (zarkon):

if 2x+4y-7z-2=0 is your equation then <2,4,7> is the normal vector

OpenStudy (zarkon):

<2,4,-7>

OpenStudy (bahrom7893):

Zarkon to the rescue!!!!

OpenStudy (istim):

Ok, got that. For the other then, it must be 4,6,-14?

OpenStudy (zarkon):

yes

OpenStudy (istim):

So from this, how would I determine the answer? Or do I need more?

OpenStudy (anonymous):

if the planes are parrallel then the cross product of their normal vectors should be 0

OpenStudy (zarkon):

two plane are parallel iff their normal vectors are parellel

OpenStudy (istim):

I'll try that cross product thing. Thanks guys. I'lll cry out if I have a problem.

OpenStudy (zarkon):

two non zero vectors \(\vec{v}\) and \(\vec{u}\) are parallel iff \(\vec{v}=c\vec{u}\) for some scalar \(c\)

OpenStudy (istim):

Pardon?

OpenStudy (zarkon):

is one of the vectors a constant multiple of the other?

OpenStudy (istim):

The first one, multiplied by 2, equals the second one.

OpenStudy (zarkon):

really?

OpenStudy (istim):

IT seemed so...

OpenStudy (zarkon):

if you have the equations 2x+4y-7z-2=0 4x+6y-14z-8=0 then you have the normal vectors <2,4,-7> and <4,6,-14> are these constant multiples of each other?

OpenStudy (istim):

That 4y was suppose to be a 3y....

OpenStudy (istim):

I did mistype that. No wonder my confusion. And probably yours...

OpenStudy (zarkon):

then <4,6,-14>=2<2,3,-7> and so they are parallel...and therefore their planes are parallel

OpenStudy (istim):

And distinct??

OpenStudy (zarkon):

yes

OpenStudy (istim):

Awesome. Thank you. This was relatively simple after all...